Find both the r-intercept and the y-intercept of the line given by this equation. 1.5x-0.7y-0.8-0 Round your answers to two decimal places x-intercept: y-intercept:
Answer #1

The x intercept is the point through which the line passes through the x axis. On this point, the value of y = 0

Substituting y = 0 in 1.5x - 0.7y - 0.8 = 0, we get 1.5x = 0.8 or x = 8/15 = 0.53

The y intercept is the point through which the line passes through the y axis. On this point, the value of x = 0

Substituting x = 0 in 1.5x - 0.7y - 0.8 = 0, we get -0.7y = 0.8 or y = -8/7 = -1.14